Solve using matrices
4x-y=5
3x-3y=0

Help please!

Answer by TimothyLamb(4379)

---
4x-y=5
3x-3y=0
---
copy and paste the above standard form linear equations in to this (matrix method) solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= 1.666666 = 1+2/3
y= 1.666666 = 1+2/3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php